# On-Chip Frequency Noise Cancellation in Nanomechanical Resonators using Cavity Optomechanics

### v1

@@ Abstract
Understanding and minimizing the sources of frequency noise in nanomechanical resonators is crucial for many sensing applications. In this work, we report an ultracoherent perimeter-mode nanomechanical resonator co-integrated with an on-chip optical cavity. This device combines low thermomechanical force noise and low detector noise, allowing us to study its intrinsic frequency fluctuations in detail. We find that the fluctuations of two mechanical modes are strongly correlated. Moreover, we demonstrate the generation of a signal at the frequency difference between the two modes directly on chip via nonlinear optomechanical transduction. This ‘difference signal’ has vastly reduced intrinsic frequency fluctuations and can be used for frequency tracking with high precision, as we establish in a proof-of-principle experiment.
@@ I Introduction
Precise tracking of mechanical eigenfrequencies is central to most applications of micro- and nanomechanical systems, including clocks, filters, and accelerometers [1, 2], as well as sensors for scanning force microscopy [3, 4, 5, 6], mass spectroscopy [7, 8], pressure sensing [9], and photothermal detection of nanoparticles [10]. In such applications, the measurement precision is limited by several sources of uncertainty [11, 12, 13, 14]: First, the thermomechanical motion of a resonator creates phase noise [11, 15] in the oscillation with a power spectral density (PSD) equal to $S_{\phi}^{\mathrm{th}}(\Omega)$, where $\Omega$ is the sideband frequency, i.e., the offset from the resonance frequency $\omega_{0}$. When measuring $\omega_{0}$, this phase noise leads to a ‘flat’ frequency noise PSD $S_{\omega}^{\mathrm{th}}(\Omega)\propto\Omega^{0}$. This noise sets a fundamental limit for the precision of frequency estimation at all time scales. Second, uncertainty in the frequency estimation is added by the readout (detection) noise. This noise leads to an apparent frequency noise $S_{\omega}^{\mathrm{det}}(\Omega)\propto\Omega^{2}$, which typically dominates the estimation of $\omega_{0}$ at short timescales. Third, coupling of the resonator to two-level systems [16] induces intrinsic frequency jitter [15], while coupling to an environment with temperature fluctuations in turn causes frequency fluctuations [17] owing to a finite thermal capacitance [11]. Together, both sources lead to a frequency noise PSD of $S_{\omega}^{\mathrm{int}}(\Omega)\propto 1/\Omega^{\alpha}$, with $\alpha$ between $0.5$ and $2$. This real frequency noise typically dominates at long timescales.
@@
Various strategies have been employed to optimize frequency stability and its estimation. The thermomechanical contribution $S_{\omega}^{\mathrm{th}}$ is minimized by designing devices with low dissipation, which was one of the central efforts in the nanomechanics community over the last decade [18, 19, 20, 21, 22, 23]. Reducing the readout noise $S_{\omega}^{\mathrm{det}}$ requires better detection sensitivity, which can be achieved with cavity optomechanics [24, 25, 26, 27, 28, 29]. By contrast, overcoming intrinsic frequency fluctuations $S_{\omega}^{\mathrm{int}}$ is more challenging, as their microscopic origin is less well understood and harder to manipulate. In addition, this contribution impacts the measurement on long timescales and is not amenable to time averaging. The influence of intrinsic frequency fluctuations can be reduced by subtracting the correlated fluctuations of various modes in post-processing [30]. However, such analysis is generally susceptible to uncorrelated detector noise and cannot always be applied to real-time measurements. Alternatively, one can drive a nonlinear resonator at an optical working point [31, 32, 33], which requires careful calibration and fine-tuning of the system. In general, it would be desirable to have a method for direct, on-chip cancellation of frequency fluctuations that requires no special calibration.
@@
In this work, we test an optomechanical platform that addresses all three frequency noise sources simultaneously. The device is an ultralow-dissipation perimeter-mode nanomechanical resonator [22] co-integrated with an on-chip optical cavity [27]. With low thermomechanical and readout noise, we can study the intrinsic frequency fluctuations over several order of magnitude in $\Omega$. Importantly, we find that the fluctuations measured on two different mechanical modes are strongly correlated. Taking advantage of the system’s nonlinear optomechanical transduction, we generate a signal at the frequency difference of the modes directly on the chip. This signal has vastly reduced frequency fluctuations and can be used as a resource for precise frequency tracking. We demonstrate this functionality in a proof-of-principle frequency measurement experiment.
II Experimental platform
@@
The optomechanical system is formed by co-integrating a polygon nanomechanical resonator [22] with a photonic crystal optical micro-cavity. The polygon resonator is suspended in the near-field of the cavity and its motion couples dispersively to the optical modes through the evanescent field. Figs. 1(a) and (b) show scanning electron microscope (SEM) micrographs of a device. The devices are fabricated out of $250\text{\,}\mathrm{nm}$ thick stoichiometric silicon nitride (Si3N4) with 1.1 $1\text{\,}\mathrm{GPa}$ deposition stress. The polygon resonator is partially thinned down to $100\text{\,}\mathrm{nm}$. The devices are then patterned using electron beam lithography and dry etching and finally released using a KOH wet process. The details of the design and fabrication of these devices are discussed in Ref. [27].
@@
Figure 1(c) illustrates our optical detection scheme. Continuous-wave laser light from a Toptica CTL 1550 is employed to drive the photonic crystal cavity at a fixed cavity detuning. We probe a cavity resonance at $1552.6\text{\,}\mathrm{nm}$ with a total linewidth of $\kappa/2\pi$ = $7.66\text{\,}\mathrm{GHz}$, see Fig. 1(d). The cavity has a free spectral range of $4.25\text{\,}\mathrm{n}\mathrm{m}$, see SI. Its DC component is used to lock the laser frequency $5.97\text{\,}\mathrm{GHz}$ red-detuned from the cavity resonance, while the AC component encodes the nanobeam’s motion. To minimize gas damping, the chip is placed in a vacuum chamber at a pressure of about $10^{-7}$ mbar. A piezo actuator, clamped to the chip, is used to excite the mechanical modes of the polygon resonator.
@@
When the laser is detuned on the side of the optical resonance, we can observe the mechanical motion through direct detection. Using a digital lock-in amplifier (Zurich Instruments MFLI), we measure the power spectral density (PSD) of the mechanical displacement driven by thermomechanical force noise, see Fig. 1(e). The peaks correspond to the thermal motion of the mechanical modes of the polygon resonator, as confirmed by FEM simulations (see insets). This PSD is calibrated in units of frequency fluctuations of the optical resonance, see SI for details. The mechanical modes shown in this frequency range correspond to the high-$Q$ fundamental out-of-plane (OOP) perimeter mode and a family of in-plane (IP) modes, also confined to the perimeter of the device. The two modes used in our experiments are the OOP mode with frequency $\omega_{1}$ = $1.070\text{\,}\mathrm{MHz}$ and the IP mode with frequency of $\omega_{2}$ = $1.092\text{\,}\mathrm{MHz}$. Using the ringdown technique, we measure quality factors of $Q_{1}=4\times 10^{7}$ and $Q_{2}=3\times 10^{5}$ for the OOP and IP modes, respectively. For these modes, we also calibrate the vacuum optomechanical coupling rates to be $g_{0,1}/2\pi=$5.2\text{\,}\mathrm{kHz}$\$ and $g_{0,2}/2\pi=$15.6\text{\,}\mathrm{kHz}$\$ (see SI for ringdown measurements and details of the calibration).
@@
We employ a digital phase-locked loop (PLL) to measure the modes’ frequencies as a function of time. The frequency noise PSD of the OOP mode shown in Fig. 1(f) is measured using different input optical powers at the same detuning from the cavity resonance, shown as a dashed grey line in Fig. 1(d). Due to the high $Q$, the thermomechanical contribution is too small to be measured and the spectra can be fit well with a model composed of only the $1/\Omega$ and detection noise contributions. We clearly see power-dependent detection noise at large sideband frequencies, while $1/\Omega$-type fluctuations at low sideband frequencies are independent of power. We use the $A/\Omega$ component of the spectra to calculate the corresponding Allan deviation as $\sigma_{\omega}=\sqrt{2A\ln{(2)}}$, where $A$ is a constant (see SI for details). These values are shown in the inset of Fig. 1(f), showing the power independence. These measurements confirm that the frequency noise for long integration times (low sideband frequencies) is dominated by $1/\Omega$-type fluctuations that are not related to optical absorption.
@@
Figure 2: Properties of signal at $\omega_{\mathrm{d}}=\omega_{2}-\omega_{1}$. Displacement PSD of demodulated around (a) $\omega_{\mathrm{d}}=\omega_{2}-\omega_{1}$ (b) $\omega_{1}$, (c) $\omega_{2}$, and (d) $\omega_{s}=\omega_{1}+\omega_{2}$. The mechanical modes at $\omega_{1}$ (red) and $\omega_{2}$ (blue) are driven with separate PLLs. (e)-(g) Spectograms of the difference signal and the displacements of the OOP and the IP modes. The color bar corresponds to the normalized PSD. (h) Double-sided frequency-PSD for $\omega_{1}$, $\omega_{2}$, and $\omega_{\mathrm{d}}$ spanning $150\text{\,}\mathrm{s}$ with a data rate of $14.3\text{\,}\mathrm{kHz}$. We use the PLLs for estimating $\omega_{1}$ and $\omega_{2}$, while $\omega_{\mathrm{d}}$ is calculated using phase-to-frequency conversion [37]. (i) Absolute Allan deviation $\sigma$ in  $\text{\,}\mathrm{Hz}$ calculated using the same traces as in (h).
III Nonlinear frequency noise cancellation
@@
Interestingly, when $\omega_{1}$ and $\omega_{2}$ are driven simultaneously, we observe optical signatures at the sum and difference frequencies ($\omega_{s,d}=\omega_{2}\pm\omega_{1}$), see Fig. 2(a)-(d). These signatures are due to the non-linear optical transduction of the mechanical signal by the cavity (see supplementary for details) that leads to mixing of the two signals [38]. More precisely, the cavity detuning is a time-varying parameter due to the mechanical motion of the resonator. This modulates the intensity of the out-coupled cavity response, where the first-order signal is at the mechanical resonance frequencies and the second-order effects are at their linear combinations. The displacement PSDs in Figs. 2(a)-(d) provide direct access to various frequency signal components. However, the $1/\Omega$-type frequency fluctuations are not easily visible in such a PSD on short timescales. To visualize the frequency fluctuations, we monitor how the PSDs change over time, see Figs. 2(e)-(g). Indeed, we find that both $\omega_{1}$ and $\omega_{2}$ change by more than $30\text{\,}\mathrm{Hz}$—hundreds of linewidths for the OOP mode—over several hours. The fluctuations of the two modes are strongly correlated but exhibit no clear pattern as a function of time, as expected for $1/\Omega$-type fluctuations. We also observe strong dependence on temperature changes (see SI). By contrast, the signal at $\omega_{\mathrm{d}}$ is significantly more stable, with a maximum change of about $0.4\text{\,}\mathrm{Hz}$, roughly two orders of magnitude smaller than that of $\omega_{1,2}$. We illustrate this in the corresponding frequency noise PSD (with uniform distribution of points along x-axis [39, 40]) in Fig. 2(h). Here, the difference mode has much lower contributions at sideband frequencies below $1\text{\,}\mathrm{Hz}$, marked by a decrease by an order of magnitude in the $1/\Omega$-type branch. One also observes how the noise floor is limited by the thermomechanical noise contribution of the mode with lower $Q$, i.e. $\omega_{2}$ to around $1.4\text{\times}{10}^{-4}\text{\,}\mathrm{Hz}^{2}\mathrm{/}\mathrm{Hz}$. In the corresponding Allan deviation $\sigma$ in Fig. 2(i), $\sigma_{\omega_{1}}$ is at its minimum around $0.9\text{\,}\mathrm{s}$ at a value of $26.1\text{\,}\mathrm{mHz}$ and starts increasing again due to thermal drifts. The difference signal, however, has a $\sigma_{\omega_{\mathrm{d}}}$ as low as $7.5\text{\,}\mathrm{mHz}$ at $\tau=$25\text{\,}\mathrm{s}$\$. For integration times beyond roughly $1\text{\,}\mathrm{s}$, the frequency fluctuations are lowered by up to a factor of 20 in the difference signal. In all of these graphs, it is evident that the frequency difference exhibits lower fluctuations than the OOP and IP modes at longer integration times (low sideband frequencies). This effect is directly tied to the origin of this ‘difference signal’ as a mixing product of modes $\omega_{1}$ and $\omega_{2}$. As the frequency fluctuations of the two real modes are correlated, their frequency difference remains approximately constant.
IV Mass sensing simulated experiment
@@
In Fig. 3, we demonstrate that the stable signal at $\omega_{\mathrm{d}}$ is potentially useful for frequency measurements. As in Fig. 2, we simultaneously drive and measure $\omega_{1}$ and $\omega_{2}$ with a PLL, while demodulating the signal at $\omega_{\mathrm{d}}$ without applied drive. Additionally, we use an active feedback loop tuned to induce a pure frequency shift $\Delta\widetilde{\omega}$ on the OOP mode at $\omega_{1}$ without affecting the mechanical linewidth (by applying a feedback in-phase with the oscillator position, resulting in softening the spring constant [41, 42, 43]). Toggling this feedback signal on and off results in frequency jumps of $\omega_{1}$ without affecting $\omega_{2}$. We refer to the frequency shift signal using the normalised gain parameter $U\in[0,1]$ where $U=1$ corresponds to the maximum $\Delta\widetilde{\omega}$ shown in the graph. This controllable frequency modulation is used to simulate the frequency jumps expected, for example, in single-particle mass sensing experiments [7, 8].
@@
In the time trace of $\omega_{1}$ in Fig. 3, it is difficult to distinguish the response to $U$ by eye, as it is partially masked by frequency fluctuations. As expected, $\omega_{2}$ does not display any systematic response to $U$ at all. For a quantitative analysis, we refer to each section between the grey dashed vertical lines as a block. We calculate the signal-to-noise ratio (SNR) as our metric, where we use the mean frequency jump size between two blocks as the signal and the standard deviation within each block to quantify the noise. We compare the SNR of the measured $\omega_{\mathrm{d}}$ against the measured $\omega_{1}$ and the calculated $\omega^{\mathrm{cal}}_{\mathrm{d}}=\omega_{2}-\omega_{1}$. For unprocessed data, we obtain an SNR of $4.1\pm 0.3$ at $\omega_{\mathrm{d}}$ in contrast with an SNR of $2.4\pm 1.7$ at $\omega_{1}$ and an SNR of $1.3\pm 0.1$ for $\omega^{\mathrm{cal}}_{\mathrm{d}}$. If we perform a simple moving-average filter over 200 data points in time (corresponding to a sync filter time of $24\text{\,}\mathrm{ms}$), the SNR at $\omega_{\mathrm{d}}$ is $6.2\pm 1$, while we obtain an SNR of $2.5\pm 1.2$ at $\omega_{1}$ and $4.5\pm 0.6$ for $\omega^{\mathrm{cal}}_{\mathrm{d}}$. The SNR of the difference signal is consistently better than that of the post-processed $\omega^{\mathrm{cal}}_{\mathrm{d}}$, as well as that of $\omega_{1}$ itself. In the SI, we present an alternative analysis based on cross-correlations, arriving at a similar result.

### v2

@@ Abstract
Understanding and minimizing the sources of frequency noise in nanomechanical resonators is crucial for many sensing applications. In this work, we report an ultracoherent perimeter-mode nanomechanical resonator co-integrated with an on-chip optical cavity. This device combines low thermomechanical force noise and low detector noise, allowing us to study its intrinsic frequency fluctuations in detail. We find that the fluctuations of two mechanical modes are strongly correlated; this holds not only for thermal drifts but also for intrinsic frequency flicker noise. Moreover, we demonstrate the generation of a signal at the frequency difference between the two modes directly on chip via nonlinear optomechanical transduction. This ‘difference signal’ has vastly reduced intrinsic frequency fluctuations and can be used for frequency tracking with high precision, as we establish in a proof-of-principle experiment.
@@
Introduction.— Precise tracking of mechanical eigenfrequencies is central to most applications of micromechanical and nanomechanical systems, including clocks, filters, and accelerometers [1, 2], as well as sensors for scanning force microscopy [3, 4, 5, 6], mass spectroscopy [7, 8], pressure sensing [9], and photothermal detection of nanoparticles [10]. In such applications, the measurement precision is limited by several sources of uncertainty [11, 12, 13, 14]: First, the thermomechanical motion of a resonator creates phase noise [11, 15] in the oscillation with a power spectral density (PSD) equal to $S_{\phi}^{\mathrm{th}}(\Omega)$, where $\Omega$ is the sideband frequency, i.e., the offset from the resonance frequency $\omega_{0}$. When $\omega_{0}$ is being measured, this phase noise leads to a ‘flat’ frequency noise PSD $S_{\omega}^{\mathrm{th}}(\Omega)\propto\Omega^{0}$. This noise sets a fundamental limit for the precision of frequency estimation at all timescales. Second, uncertainty in the frequency estimation is added by the readout (detection) noise. This noise leads to an apparent frequency noise $S_{\omega}^{\mathrm{det}}(\Omega)\propto\Omega^{2}$, which typically dominates the estimation of $\omega_{0}$ at short times. Third, coupling of the resonator to two-level systems [16, 17] induces intrinsic frequency flicker, also known as jitter [15, 18], whereas coupling to an environment with temperature changes causes frequency drifts and fluctuations [19], owing to a finite thermal capacitance [11]. Together, both sources lead to a frequency noise PSD $S_{\omega}^{\mathrm{int}}(\Omega)\propto 1/\Omega^{\alpha}$, with $\alpha$ between $0.5$ and $2$. This real frequency noise typically dominates at long times.
@@
Various strategies have been used to optimize frequency stability and its estimation. The thermomechanical contribution $S_{\omega}^{\mathrm{th}}$ is minimized by designing devices with low dissipation, which was one of the central efforts in the nanomechanics community over the last decade [20, 21, 22, 23, 24, 25]. Reducing the readout noise $S_{\omega}^{\mathrm{det}}$ requires greater detection sensitivity, which can be achieved with cavity optomechanics [26, 27, 28, 29, 30, 31]. By contrast, overcoming intrinsic frequency fluctuations $S_{\omega}^{\mathrm{int}}$ is more challenging, as their microscopic origin is less well understood and harder to manipulate. In addition, this contribution is not amenable to time averaging as it impacts the measurement on long timescales. The influence of intrinsic frequency fluctuations can be reduced by subtracting the correlated fluctuations of various modes in postprocessing [32]. However, such analysis is generally susceptible to uncorrelated detector noise and cannot always be applied to real-time measurements. Alternatively, one can drive a nonlinear resonator at an optical working point [33, 34, 35, 36], which requires careful calibration and fine-tuning of the system. Until now, no method has been shown that offers on-chip, calibration-free cancellation of frequency fluctuations.
@@
In this work, we address this challenge with an on-chip optomechanical platform. The device is an ultralow-dissipation perimeter-mode nanomechanical resonator [24] co-integrated with an on-chip optical cavity [29]. With low thermomechanical and readout noise, we can study the intrinsic frequency fluctuations over several orders of magnitude in $\Omega$. Importantly, we find that the fluctuations measured on two different mechanical modes are strongly correlated. Taking advantage of the system’s nonlinear optomechanical transduction, we generate a signal at the frequency difference of the modes directly on the chip. This signal has vastly reduced frequency fluctuations and can be used as a resource for precise frequency tracking. We demonstrate this functionality in a proof-of-principle frequency measurement experiment.
@@
Experimental platform.— The optomechanical system is formed by co-integration of a polygon nanomechanical resonator [24] with a photonic crystal optical micro-cavity. The polygon resonator is suspended in the near field of the cavity, and its motion couples dispersively to the optical modes through the evanescent field. Figs. 1(a) and (b) show scanning electron microscope micrographs of a device. The devices are fabricated from $250\text{\,}\mathrm{nm}$-thick stoichiometric silicon nitride (Si3N4) with 1.1 $1\text{\,}\mathrm{GPa}$ deposition stress. The polygon resonator is partially thinned down to $100\text{\,}\mathrm{nm}$. The devices are then patterned by electron beam lithography and dry etching, and are finally released with the use of a KOH wet process. The details of the design and fabrication of these devices are discussed in Ref. [29].
@@
Figure 1(c) illustrates our optical detection scheme. Continuous-wave laser light from a Toptica CTL 1550 laser is to drive the photonic crystal cavity at a fixed cavity detuning. We probe a cavity resonance at $1552.6\text{\,}\mathrm{nm}$ with total linewidth $\kappa/2\pi$ = $7.66\text{\,}\mathrm{GHz}$; see Fig. 1(d). The cavity has a free spectral range of $4.25\text{\,}\mathrm{n}\mathrm{m}$ (see Supplemental Material A for the data and calibration [40]). Its dc component is used to lock the laser frequency $5.97\text{\,}\mathrm{GHz}$ red-detuned from the cavity resonance, while the ac component encodes the nanobeam’s motion. To minimize gas damping, the chip is placed in a vacuum chamber at a pressure of about $1\text{\times}{10}^{-7}\text{\,}\mathrm{mbar}$. A piezoelectric actuator, clamped to the chip, is used to excite the mechanical modes of the polygon resonator.
@@
We detune the laser by $\Delta=$5.97\text{\,}\mathrm{GHz}$\$ to the red side of the resonance and observe mechanical motion through direct detection. Using a digital lock-in amplifier (Zurich Instruments MFLI), we measure the PSD of the mechanical displacement driven by thermomechanical force noise, see Fig. 1(e). The peaks correspond to the thermal motion of the mechanical modes of the polygon resonator, as confirmed by FEM simulations (see insets). This PSD is calibrated in units of frequency fluctuations of the optical resonance, see Supplemental Material D for details. The mechanical modes shown in this frequency range correspond to the high-$Q$ fundamental out-of-plane (OOP) perimeter mode and a family of in-plane (IP) modes, also confined to the perimeter of the device. The two modes used in our experiments are the OOP mode with frequency $\omega_{1}$ = $1.070\text{\,}\mathrm{MHz}$ and the IP mode with frequency of $\omega_{2}$ = $1.092\text{\,}\mathrm{MHz}$. Using the ringdown technique, we measure quality factors of $Q_{1}=4\times 10^{7}$ and $Q_{2}=3\times 10^{5}$ for the OOP and IP modes, respectively. For these modes, we also calibrate the vacuum optomechanical coupling rates to be $g_{0,1}/2\pi=$5.2\text{\,}\mathrm{kHz}$\$ and $g_{0,2}/2\pi=$15.6\text{\,}\mathrm{kHz}$\$ (see Supplemental Material 5 for ringdown measurements and details of the calibration).
@@
We use a digital phase-locked loop (PLL) to measure the modes’ frequencies as a function of time. The frequency noise PSD of the OOP mode shown in Fig. 1(f) is measured for different input optical powers at the same detuning from the cavity resonance, shown as a dashed gray line in Fig. 1(d). Because of the high $Q$, the thermomechanical contribution is too small to be measured, and the spectra can be fit well with a model composed of only the $1/\Omega$-component shown as a dashed gray line in Fig. 1(f) in addition to detection noise contributions. We clearly see power-dependent detection noise at large sideband frequencies, while $1/\Omega$-type fluctuations at low sideband frequencies are independent of power. Independence of these fluctuations from mechanical drive power, pressure [41] and temperature fluctuations is also shown in Supplemental Material E. We use the $A/\Omega$ component of the spectra to calculate the corresponding Allan deviation for long integration times as $\sigma_{\omega}=\sqrt{2A\ln{(2)}}$, where $A$ is a constant (see Supplemental Material C for details). These values are shown in the inset in Fig. 1(f), showing the power independence. These measurements confirm that the frequency noise for long integration times (low sideband frequencies) is dominated by $1/\Omega$-type fluctuations that are not related to optical absorption.
@@
Figure 2: Properties of signal at $\omega_{\mathrm{d}}=\omega_{2}-\omega_{1}$. Displacement PSD of demodulated signal around (a) $\omega_{\mathrm{d}}=\omega_{2}-\omega_{1}$ (b) $\omega_{1}$, (c) $\omega_{2}$, and (d) $\omega_{s}=\omega_{1}+\omega_{2}$. The mechanical modes at $\omega_{1}$ (red) and $\omega_{2}$ (blue) are driven with separate PLLs. (e)-(g) Spectograms of the difference signal and the displacements of the OOP and the IP modes. The colour bar corresponds to the normalized PSD. (h) Double-sided frequency-PSD for $\omega_{1}$, $\omega_{2}$, and $\omega_{\mathrm{d}}$ spanning $150\text{\,}\mathrm{s}$ with a data rate of $14.3\text{\,}\mathrm{kHz}$. We use the PLLs for estimating $\omega_{1}$ and $\omega_{2}$, while $\omega_{\mathrm{d}}$ is calculated using phase-to-frequency conversion [42]. The dashed-gray line is the fit for the $1/\Omega$-component of the frequency noise PSD. (i) Absolute Allan deviation $\sigma$ in  $\text{\,}\mathrm{Hz}$ calculated using the same traces as in (h).
@@
Nonlinear frequency noise cancellation.— Interestingly, when $\omega_{1}$ and $\omega_{2}$ are driven simultaneously, we observe optical signatures at the sum and difference frequencies ($\omega_{d}=\omega_{2}-\omega_{1}$, $\omega_{s}=\omega_{2}+\omega_{1}$), see Fig. 2(a)-(d). These signatures are due to the nonlinear optical transduction of the mechanical signal by the cavity (see Supplemental Material F for details) that leads to mixing of the two signals [43]. More precisely, the cavity detuning is a time-varying parameter due to the mechanical motion of the resonator. This detuning modulates the intensity of the out-coupled cavity response, where the first-order signal is at the mechanical resonance frequencies, and the second-order effects are at their linear combinations. The displacement PSDs in Figs. 2(a)-(d) provide direct access to various frequency signal components. However, the $1/\Omega$-type frequency fluctuations are not easily visible in such a PSD on short timescales. To visualize the frequency fluctuations, we monitor how the voltage PSDs change over time, see Figs. 2(e)-(g). Indeed, we find that both $\omega_{1}$ and $\omega_{2}$ change by more than $30\text{\,}\mathrm{Hz}$—hundreds of linewidths for the OOP mode—over several hours. The fluctuations of the two modes are strongly correlated but exhibit no clear pattern as a function of time, as expected for $1/\Omega$-type fluctuations. We also observe strong dependence on temperature changes (see Supplemental Material E.1,[44]). By contrast, the signal at $\omega_{\mathrm{d}}$ is significantly more stable, with a maximum change of about $0.4\text{\,}\mathrm{Hz}$, roughly two orders of magnitude smaller than that of $\omega_{1,2}$.
We illustrate the improvement in the corresponding frequency noise PSD (with uniform distribution of points along the x-axis [45, 46]) in Fig. 2(h). The two modes at $\omega_{1}$ and $\omega_{2}$ follow the same $1/\Omega$ slope, shown as a dashed gray line. The difference mode, however, has much lower contributions at sideband frequencies below $1\text{\,}\mathrm{Hz}$, marked by a decrease by an order of magnitude in the $1/\Omega$-type branch. We observe that the noise floor is limited by the thermomechanical noise contribution of the mode with lower $Q$, i.e. $\omega_{2}$, to around $1.4\text{\times}{10}^{-4}\text{\,}\mathrm{Hz}^{2}\mathrm{/}\mathrm{Hz}$. In the corresponding Allan deviation $\sigma$ in Fig. 2(i), $\sigma_{\omega_{1}}$ is at its minimum value $26.15(0.03)\text{\,}\mathrm{mHz}$ around $0.9\text{\,}\mathrm{s}$ and starts increasing again for longer times. The difference signal, however, has a $\sigma_{\omega_{\mathrm{d}}}$ up to a factor of 20 lower than for the OOP and IP modes at long integration times, reaching $7.54(0.01)\text{\,}\mathrm{mHz}$ at $\tau=$25\text{\,}\mathrm{s}$\$. This effect is directly tied to the origin of this ‘difference signal’ as a mixing product of modes $\omega_{1}$ and $\omega_{2}$. As the frequency fluctuations of the two real modes are correlated, their frequency difference remains approximately constant. As we show in Appendix D in the Supplementary Material D, the correlated fluctuations stem both from environmental temperature changes and intrinsic flicker noise, and are both reduced by our optomechanical mixing technique. Both sources can be identified separately and can be shown to follow different power laws for frequencies beyond $0.01\text{\,}\mathrm{Hz}$.
@@
Feedback frequency shift experiment.— In Fig. On-Chip Frequency Noise Cancellation in Nanomechanical Resonators using Cavity Optomechanics, we demonstrate that the stable signal at $\omega_{\mathrm{d}}$ is potentially useful for frequency measurements. As in Fig. 2, we simultaneously drive and measure $\omega_{1}$ and $\omega_{2}$ with a PLL, while demodulating the signal at $\omega_{\mathrm{d}}$ without applied drive. Additionally, we use an active feedback loop tuned to induce a pure frequency shift $\Delta\widetilde{\omega}$ on the OOP mode at $\omega_{1}$ without affecting the mechanical linewidth (by applying a feedback in phase with the oscillator position, resulting in softening of the spring constant [47, 48, 49]). Toggling this feedback signal on and off results in frequency jumps of $\omega_{1}$ without affecting $\omega_{2}$. We refer to the frequency shift signal using the normalized gain parameter $U\in[0,1]$ where $U=1$ corresponds to the maximum $\Delta\widetilde{\omega}$ shown in the graph. This controllable frequency modulation is used to simulate the frequency jumps expected, for example, in single-particle mass sensing experiments [7, 8].
Figure 3. Measuring frequency jumps with high precision at $\omega_{\mathrm{d}}$. (a) Effect of frequency jumps $\Delta\omega$ acting on $\omega_{1}$, but not on $\omega_{2}$. (b) Expected frequency jumps $\Delta\widetilde{\omega}$ when a dispersive feedback on mode $\omega_{1}$ is toggled on and off. The two modes at $\omega_{1}$ and $\omega_{2}$ are driven by separate PLLs. Frequency shifts (c) $\Delta\omega_{1}$, (d) $\Delta\omega_{2}$, and (e) $\Delta\omega_{\mathrm{d}}$ are simultaneously measured as a function of time.
@@
In the time trace of $\omega_{1}$ in Fig. On-Chip Frequency Noise Cancellation in Nanomechanical Resonators using Cavity Optomechanics, it is difficult to distinguish the response to $U$ by eye, as it is partially masked by frequency fluctuations. As expected, $\omega_{2}$ does not display any systematic response to $U$ at all. For a quantitative analysis, we refer to each section between the vertical gray dashed lines as a block. We calculate the signal-to-noise ratio (SNR) as our metric, where we use the mean frequency jump size between two blocks as the signal and the standard deviation within each block to quantify the noise. We compare the SNR of the measured $\omega_{\mathrm{d}}$ against the measured $\omega_{1}$ and the value $\omega^{\mathrm{cal}}_{\mathrm{d}}=\omega_{2}-\omega_{1}$ calculated in postprocessing [32]. For unprocessed data, we obtain an SNR of $4.1\pm 0.3$ at $\omega_{\mathrm{d}}$ in contrast with an SNR of $2.4\pm 1.7$ at $\omega_{1}$ and an SNR of $1.3\pm 0.1$ for $\omega^{\mathrm{cal}}_{\mathrm{d}}$, noting an increase by a factor of 3.2 at the same frequency. When we perform a simple moving-average filter over 200 data points in time (corresponding to a sync filter time of $24\text{\,}\mathrm{ms}$), the SNR at $\omega_{\mathrm{d}}$ is $6.2\pm 1$, while we obtain an SNR of $2.5\pm 1.2$ at $\omega_{1}$ and $4.5\pm 0.6$ for $\omega^{\mathrm{cal}}_{\mathrm{d}}$. The SNR of the difference signal is consistently better than that of the post-processed $\omega^{\mathrm{cal}}_{\mathrm{d}}$, as well as that of $\omega_{1}$ itself. In the SI, we present an alternative analysis based on cross-correlations, arriving at a similar result.